A) The spot price of the British pound is currently $2.00. If the risk-free interest rate on 1-year Government bonds is 4% in the United States and 6% in the United Kingdom,

what must be the forward price of the pound for delivery 1 year from now?

B) Assume that the spot price of gold is $1,500 per troy ounce, the risk-free interest rate is 2%, and

storage and insurance costs are zero.

1) What should be the forward price of gold for delivery in 1 year?

2) If the futures price is $1550, develop a strategy that can bring risk-free arbitrage profits.

3) Calculate the profit that you can make by following that arbitrage strategy.
Answer #1

Forward price = Spot price(1+Interest rate US)/(1+Interest Rate UK)

= 2*(1+4%)/(1+6%)

= $1.9623/Pound

B) Forward price = Spot rate*(1+Risk free rate)

= 1500*1.02

= $1,530

2)Take loan of $1500 and buy today at 1500 sell forward at $1550

Repay Loan = 1500*1.02 = $1530

3)Arbitrage Profit = $20